Project

General

Profile

NDOS Commissioning Meeting 2011/10/20

Attendance: Jon, Gary, Mat, Nick, Minerba, Athans, Martin, Gavin

Agenda:

  1. APD Redistribution Update
  2. DCS Update
  3. DAQ Update
  4. AOB

APD Redistribution Update

  • Mat: two more batches have been moved this week:
    • one batch to MC horizontals, only 5/11 were successful. Two appear strange, perhaps a problem with FEBs
    • yesterday, 12 were moved, 11/12 were successful.
    • ignoring problems with MC, success rate is still ~80%
  • Jon/Mat: by shifting tracking planes by one or two, we can reduce # of APDs to move by 6 by using existing APDs in diblock 3. We will just go ahead and do that.
  • two DCMs were removed from the NDOS for repair, they have been re-installed and we have gained back two FEBs.

DCS Update (Martin)

  • cooling enabled mask was updated (
  • had a problem with one of their machines, but it has been recovered.

DAQ Update

  • nada

AOB

  • nada